# Sorting-based FPGA Sliding Window Aggregation Engine without off-chip Memories

### v2

@@ Abstract.
Aggregation queries are a series of computationally-demanding analytics operations on grouped and time series data. They include tasks such as summation or finding the median among the items of a group sharing a group ID, and within a specified number of the last observed tuples for sliding window aggregation (SWAG). They have a wide range of applications including in database analytics, operating systems, bank security and medical sensors. Existing challenges include the hardware complexity that comes with efficiently handling per-group states using hash-based approaches. This paper presents a pipelined and adaptable approach for calculating a wide range of aggregation queries with high throughput. It is then adapted for SWAG to achieve up to 476x speedup over the CPU of the same platform. It outperforms the state-of-the-art such as by being able to process 7.14x more tuples per second, and support 4x the window sizes with a fraction of the resources and no DRAM.
@@
FPGA, reconfigurable, group-by-aggregate, engine, prefix scan, sliding window, SWAG, key-value pairs
1. Introduction
@@
Aggregation is the task of combining raw data to produce more useful and summarised knowledge. It appears in many disciplines, especially relating to distributed and parallel programming models, such as MapReduce (Tangwongsan et al., 2018), where multiple workers/nodes produce a partial result that then needs aggregating. The worker notion is also popular in FPGA designs (Owaida et al., 2017). Equivalently, an established software design pattern is the fork-join model, which is a prominent concurrency pattern with processes or threads calculating partial results (De Wael et al., 2014).
In data science, there are numerous operations that can contribute to the usefulness of these computational models. One way is with the column and key-value-store paradigms, which are also increasingly adopted in high-performance analytics databases, as opposed to row-store. This is because of the potential to stream whole columns, which increases data localities and linearises the memory access pattern for certain analytics operations, also benefiting FPGA acceleration.
@@
One common aggregation application is the group-by aggregate operator for relational databases. It provides a summary of rows belonging to (having) the same group ID. An example SQL (structured query language) snippet is provided in algorithm 1, where a sum operation is applied to all keys (table0.key2) of the same group (table0.key1). Other than summation, SQL aggregate functions include minimum, maximum, average and count. These are sometimes combined with more advanced options such as for calculating the distinct key count per group ID.
A similar category of queries is sliding window aggregation (SWAG) (Tangwongsan et al., 2018). A time domain is associated with the data, as with queries specifying a time range to form windows for aggregation. Due to the fact that the window formation is applied on-the-fly, near where the data is gathered, the stream is considered partly sorted with regard to accompanying timestamps or the implied order of the data. SWAG usually yields a random memory access pattern to calculate, as with fine-tuned CPU algorithms (Theodorakis et al., 2018). An example is when finding the median, as even with FPGAs, current designs require caching of larger states (Geethakumari and Sourdis, 2023). This complicates the design by increasing the on-chip and off-chip memory requirements for supporting/approaching the worst case (as with multiple inputs hashing to the same slot for hashsets).
@@
This paper presents a novel design for a group-by-aggregate engine that integrates computation inside high-throughput and efficient structures such as the prefix scan. This engine is then adapted for performing SWAG queries, by adding additional functionality to an accompanying sorter. The proposed SWAG engine is highly memory-efficient to the point that it achieves equivalent functionality to highly complex competing designs without their main requirement for an off-chip DRAM (Ramakrishnan Geethakumari et al., 2019; Geethakumari and Sourdis, 2021a, b, 2023) or compression (Geethakumari and Sourdis, 2023). The internal state fits inside on-chip memories such as block-RAM. The main idea behind this efficiency is to replace the need for large indexed or hashed states by offloading the design complexity to small-scale sorting, as it eliminates any fragmentation that would be caused by the need to maintain per-group information.
@@
  • A state-of-the-art high-throughput DRAM-less sliding window aggregation (SWAG) engine based on the group-by-aggregate engine.
@@
  • A series of modifications for a high-throughput sorter to be able to append group cardinality to each tuple for the purposes of supporting the SWAG median operator.
@@
  • System-based and out-of-context evaluation of the proposed designs.
2. Background
An emerging need in data analytics is the use of programming models and related infrastructure to be able to benefit from the increasingly-wide memories and datapaths (Papaphilippou et al., 2023). A more traditional alternative is to have multiple workers and ports, as with FPGA-based database designs with multi-threading-equivalent arrangements. However, since the memory access pattern is highly-affected by this design choice, the hardware complexity is also affected, such as by co-designing cache or hash structures on the programmable logic.
Figure 1. Modular design of the distinct count query (right) with example data (Papaphilippou et al., 2020) in orange, based on the PRRA implementation (left) (Papaphilippou et al., 2023).
One recent example of a throughput-optimised building block is the reverse butterfly-based “parallel round-robin arbiter” (PRRA (Papaphilippou et al., 2023)), which removes sparsity from high-throughput data. It is essentially a round-robin load balancer, but it retains the order of the data to facilitate computation. Figure 1 illustrates how a group-by aggregation task (distinct count per group ID) can be built in a modular way using a PRRA. One of the aims of the presented work is further simplification of such pipelines for group-by aggregation, as we show that they have certain amounts of logic redundancy (as with the implementation discussion of section 4.5).
Figure 2. SWAG example for the minimum operator.
The main target of the presented work is sliding-window aggregation (SWAG). This family of queries operate on streaming data, and perform a filtering step that evaluates a statement relating to time to form windows for aggregation. The input mainly data consists of (group, key) tuples, also namely key-value pairs. The time is stated either by the natural order of their appearance in the input, or by appended timestamps (Oge et al., 2013). The scope of our work only regards the time being represented by the order of the tuples as they arrive. An example SWAG query is usually of the form “calculate the median of the last WS elements per group ID every WA tuples” (Geethakumari and Sourdis, 2023), where median can be replaced by any supported aggregate operator, WS is the window size, and WA is the distance between the sampled windows. Figure 2 shows an example for the minimum operator, where on every 2 tuples (WA) we find the minimum key per group within the last 4 tuples (WS).
3. Proposed solution
Figure 3. Five-step pipeline of the proposed group-by-aggregate engine.
In this section, we first propose a novel design for performing a common analytics set of queries (group-by-aggregate, section 3.1). This engine is then adapted to perform sliding window aggregation queries (SWAG, section 3.2). In order to support applicability to a wider range of query operators including median, an accompanying sorter is extended with group-cardinality functionality (section 3.3). The main theme of these designs is high-throughput, as they can accept $P$ elements per FPGA cycle, where $P$ is the degree of parallelism, aiming at scalability for saturating wide datapaths.
3.1. Group-by-aggregate engine
The proposed group-by-aggregate engine is a high-throughput pipeline for calculating a wide range of group-by-aggregate queries. It also works as a non-blocking building block, as it never asserts backpressure on the stream (i.e. the “ready” signal is high). Based on the query classification from (Tangwongsan et al., 2018), the supported types are Sum-like, Max-like and Sketch-like.
Figure 3 introduces the architecture of the proposed group-by-aggregate engine. A rolling prefix scan is used to calculate the permutation indices for the permutation network (reverse butterfly (Papaphilippou et al., 2023)) to achieve stream compaction. Here, the engine exploits the prefix-scan topology to calculate and append aggregation information to the data, before filtering redundant tuples (already aggregated). The engine expects a sorted stream consisting of ($group_{i},key_{i}$) tuples, arriving in batches of size $P$ per cycle. The input is assumed dense, as the engine either receives $P$ or 0 tuples in a cycle, according to other latencies at the system-level. In its output, the aggregation engine can create sparsity, hence the usefulness of the compaction functionality of the permutation network.
Steps (a) and (b)
@@ 3.1. Group-by-aggregate engine
The stage (a) in the pipeline is responsible for buffering up to one extra batch of size $P$, so that the entities (nodes) $t$ are able to (b) distinguish and mark each last tuple per $group$. Essentially, this marking is done by setting their $last$ bit high, representing the last tuple for each group. These last tuples are important, because they will be those that will carry the aggregate results, as they will be able to “see” all previous values of their $group$ before being extracted. Thus, these bits will also be used as valid-bits for the reverse butterfly network, to denote a valid tuple, also containing the summarised information after the end of the pipeline.
Step (c)
@@
The entities named $n$ are distributed inside the adapted prefix scan to calculate the permutation index of the tuples to achieve the round-robin effect by the reverse butterfly network, as with the PRRA (Papaphilippou et al., 2023). At the same time they also operate on the $key$ fields, which later becomes the $result$ field containing the aggregation summary per group. For instance, if the query function is to find the sum of all the keys for each of the group IDs, the entities $n$ not only perform a cumulative sum for the reverse butterfly $index$, but also for the key sum per group, until the last entry of the same group is processed to carry the full sum of the group. An equality check on the group field with the previous entry ($group_{i}\stackrel{{\scriptstyle?}}{{=}}group_{i-1}$) is enough for knowing where not to propagate the previous result. Using sum as an example, the start of a new $group$ would leave the current $key$ intact instead of adding the result of the last tuple.
@@
In order to also support the distinct count operator (as in figure 1), the entities $n$ also need to follow a simple distributed algorithm: if the propagated minimum from the input node of the larger key differs from the smaller key, then the distinct counts are coming from disjoint sets, hence $n$ adopts their sum. If the minimum is equal to the smaller key of the two, then $n$ adopts this sum of distinct counts, but subtracts 1 for taking care of the double counting of the common key.
Step (d)
@@
The entities $n^{\prime}$ are special versions of the entities $n$ and their main functionality is to output the originally-marked tuples as valid to the permutation network, while replacing their key field with the result. The main difference with $n$ is that sometimes the partial result information cannot be contained in the $key$ field, and they are the entrypoint for taking additional signals into account for the final result.
@@
For instance, if $function\_select$ indicates that the operation is to find the average values per group, the entities $n$ operate the same as in the sum operation. However, it is the entity $n^{\prime}$ that will divide the result by the corresponding group tuple count (cardinality). In such a case, the entities $n$ also include internal signals for counting the number of tuples in the group in a similar fashion to the sum operation, but adding 1 instead of the key. A related difference is with the count signals of $n^{\prime}$, since $n^{\prime}$ are responsible for the rolling effect, the propagated count of the bottom $n^{\prime}$ is able to count beyond $P$ elements, e.g. by having a 32-bit width. All $n^{\prime}$ elements use this value plus the $(log_{2}P+1)$-bit wide count from the prefix sum to provide the final 32-bit count in the result, when it is requested such as from a count query. According to the complexity of the supported operations, additional signals may be carried in the network, such as the minimum, that becomes of use when calculating the distinct count per group ID.
Step (e)
@@
Finally, the reverse butterfly network permutes the batch of results to achieve the round-robin effect in parallel, and is inline with (Papaphilippou et al., 2023). The reverse butterfly network is used as an optimal-complexity switch, though other permutation networks can also be used. The $last$ flag is now renamed to a $valid$ bit, as it now indicates that a port has valid result.
@@
Figure 4. Colour-coded numerical example: finding the average per group.
Figure 5. High-throughput sliding window aggregation. The main functionality is to select the min., median and max. values per group ID.
@@
Figure 4 introduces a numerical example on how a single batch of $P$ tuples is processed through the pipeline for $P=4$. Initially, the 4 tuples have group IDs ranging from {a,b,c}, and they are sorted. The last tuple for each group is known for at least the first 3, plus the remaining one because of step $(a)$. In this example, it is easy to see that (a,2) and (b,4) are marked with T for true as the last elements with that group ID (the example assumes that the next batch also starts with c, hence (c,9) becomes (c,9,F) on the next pipeline stage. Then, the marked tuples pass through the prefix scan, which appends to them (incrementally): a cumulative sum on their values (to facilitate sum-related operations), a cumulative sum on their valid bits (to calculate the number of tuples per group), and a cumulative sum on their recently-marked last bits (that will be used as an index to the reverse butterfly network for the round-robin effect). The first two sums are in the scope of their corresponding group IDs, and it is achieved by simply comparing their group IDs before interacting with each other according to the prefix scan topology. The next step is the additional stage in the prefix scan, where the $n^{\prime}$ entities finalise the results and also update the counts according to the previous batch. Here, (a,2) and (b,4), being selected as the tuples to carry the aggregated information for their groups a and b respectively, they replace their keys with the average values per group (e.g. by dividing 12 by 2 for b). The permutation indices are not altered in this case, remaining at 0 and 1 respectively, assuming this was the first batch, or that the previous batch has outputted its last aggregate result on index $P-1$.
3.2. Sliding window aggregation engine
The design from section 3.1 is adapted to perform sliding-window aggregation (SWAG) queries. The main difference from the non-SWAG engine is that the input is bounded by windows. The context of the input has a fixed user-defined window size (WS), and a simple buffering mechanism provides these windows, while skipping by window advance (WA), as explained in the background (section 2). Additionally, instead of relying on an external sorter, a sorter with a special functionality needs to be part of the SWAG engine.
Step (a)
[…]
Step (b)

### v3

# Efficient Adaptable High-throughput Streaming Aggregation Engines

@@ Abstract
Aggregation queries are a series of computationally-demanding analytics operations on counted, grouped or time series data. They include tasks such as summation or finding the median among the items of the same group, and within a specified number of the last observed tuples for sliding window aggregation (SWAG). They have a wide range of applications including database analytics, operating systems, bank security and medical sensors. Existing challenges include the hardware complexity that comes with efficiently handling per-group states using hash-based approaches. This paper presents an adaptable pipeline for calculating a wide range of aggregation queries with high throughput. It is then adapted for SWAG and achieves up to 476x speedup over the CPU core of the same platform. It achieves unparalleled levels of performance and functionality such as a throughput of 1 GT/s on our setup for SWAG without groups, and more advanced operators with up to 4x the window sizes than the state-of-the-art with groups as an approximation for SWAG featuring per-group windows using a fraction of the resources and no DRAM.
Index Terms:
@@
FPGA, reconfigurable, group-by-aggregate, engine, prefix scan, sliding window, SWAG, key-value pairs, median
I Introduction
@@
Aggregation is the task of combining raw data to produce more useful and summarised knowledge. It appears in many disciplines, especially relating to distributed and parallel programming models, such as MapReduce [1], where multiple workers/nodes produce a partial result that needs aggregating. The worker notion is also popular in FPGA designs [2]. Equivalently, an established software design pattern is the fork-join model, which is a prominent concurrency pattern with processes or threads calculating partial results [3].
An emerging need in data analytics is the use of programming models and related infrastructure to be able to benefit from the increasingly-wide memories and datapaths [4]. A more traditional alternative is to have multiple workers and ports, as with FPGA-based database designs with multi-threading-equivalent arrangements. However, since the memory access pattern is highly-affected by this design choice, the hardware complexity is also affected, such as by co-designing cache or hash structures on the programmable logic. In the software world, this has already been reflected in the prominence of the key-value store paradigm, where the simplified storage model can enable higher throughput [5].
@@
One typical aggregation application is the group-by aggregate operator for relational databases. It provides a summary of rows belonging to (having) the same group ID. An example SQL (structured query language) snippet is provided in algorithm 1, where a sum operation is applied to all keys (table0.key2) of the same group (table0.key1). Other than summation, SQL aggregate functions include minimum, maximum, average and count. These are sometimes combined with more advanced options such as for calculating the distinct key count per group ID.
@@
A similar category of queries is sliding window aggregation (SWAG) [1]. A time domain is associated with the data, as with queries specifying a time range to form windows for aggregation. Due to the fact that the window formation is applied on-the-fly, near where the data is gathered, the stream is considered partly sorted with regard to accompanying timestamps or the implied order of the data. SWAG usually yields a random memory access pattern to calculate, including for fine-tuned CPU algorithms [6]. An example is when finding the median, as even with FPGAs, current designs require caching of larger states [7]. This complicates the design by increasing the on-chip and off-chip memory requirements for supporting/approaching the worst case (as with multiple inputs hashing to the same slot for hashsets).
Figure 1: Five-step pipeline of the proposed aggregation pipeline.
This paper presents a novel design for aggregation that integrates computation inside scalable structures such as the prefix scan and the butterfly network that process multiple elements per-cycle (high-throughput). The idea behind this design is the exploitation of those structures originally used for highly-optimised compaction [4] to simultaneously append aggregation information, leading to minimal resources (efficiency). The resulting base pipeline supports a plethora of operators and use cases (adaptable), as with the proposed adaptations for performing SWAG queries, by accompanying an on-chip sorter-cardinality counter. The resulting SWAG engine is highly memory-efficient to the point that it achieves similar functionality to highly-complex competing designs without their limiting requirement for an off-chip DRAM or compression. Conveniently, all logical states fit entirely inside on-chip memories such as block RAM (BRAM). For applications such as SWAG, the concept behind its advantageous system behaviour is the replacement of the need for large indexed or hashed structures by offloading the design complexity to sorting, as it eliminates fragmentation originally caused by maintaining per-group information. FPGA-based high-throughput sorters are thoroughly studied in literature [8, 9, 10, 11], but in order to seamlessly support the median operator in our SWAG engine, innovative lightweight algorithmic modifications are introduced.
@@
  • A state-of-the-art DRAM-less sliding window aggregation engine optimised for the single-window use case.
@@
  • A series of modifications for a high-throughput sorter so that it can append group cardinality to each tuple for the purposes of supporting the SWAG median operator.
@@
  • Open-source system-based and out-of-context evaluation of the proposed designs.
This journal paper extends the similarly-named conference poster “Efficient Adaptable Streaming Aggregation Engine” [12] especially with adaptations in novel sliding-window aggregation engines (section III) and a comprehensive theoretical and system-level evaluation (section IV).
II Aggregation engine
In this section, we propose a novel pipeline as a building block for performing a range of commonly used analytics operators. This engine is later adapted to perform a wider selection of queries with sliding windows (section III-B).
The main idea behind the design is that the pre-existing prefix-scan topology of a compaction circuit can be exploited and reused to also calculate and append aggregation information to the data, before filtering redundant tuples. The focus of this pipeline is high-throughput, as it is always able to process up to $P$ elements per cycle. It is inspired by the PRRA architecture [4], where a rolling prefix scan is used to calculate the permutation indices for the permutation network (reverse butterfly) to achieve stream compaction (see figure 16, left).
II-A Supported functionality
The accepted input stream can include a group associated with each value (key-value pairs), or without, as with the simpler task of summarising plain integers. The engine expects a sorted stream consisting of ($group_{i},key_{i}$) tuples, arriving in batches of size $P$ per cycle. $P$ can be adjusted at synthesis-time to accommodate wide datapaths accordingly (elaborated in section IV-G). The input is assumed dense, since the engine either receives $P$ or 0 tuples in a cycle, according to other latencies at the system-level. The $group_{i}$ part is optional, as it can be configured to accept number streams instead, which is equivalent to associating all values to the same group.
In terms of operators, due to the prefix scan component, the pipeline supports (a superset of) associative operators, within the context of the values belonging to the same group. That is $a\oplus(b\oplus c)=(a\oplus b)\oplus c\ \forall\ a,b,c\in S$, where $S$ is the value set, such as all possible 32-bit integers. The studied operators include minimum, maximum and summation, count. It also supports combinations of the above, such as the average, by simultaneously calculating summation and count, and then combining the results near its output. When the input is entirely sorted including by-the-value, this extends its applicability, such as for performing distinct count (essentially an associative operation when the input is sorted, see step $(c)$).
This design is a non-blocking building block working as a pipeline, i.e. it never asserts backpressure on the stream, as with the PRRA architecture. On the contrary, aggregation implies summarisation, hence the need to remove a newly generated intermediate sparsity with the compaction functionality of the permutation network [4]. For most aggregation tasks, this will result in only the last tuple per group being considered useful, since it is repurposed to carry all the relevant accumulated information, while the remaining aforementioned “redundant” tuples may carry partial results.
II-B Pipeline steps
Figure 1 introduces the architecture of the proposed aggregation design, a high-throughput pipeline for calculating a wide range of aggregation tasks including group-by.
Step (a) Buffering
this step is responsible for buffering up to one extra batch of size $P$, so that the entities (nodes) $t$ of the next step (b) are able to distinguish and mark each last tuple per groupgroup.
Step (b) Marking last keys
@@ II-B Pipeline steps
essentially, this marking is done by setting the tuple’s lastlast bit high, representing the last tuple for each group. These last tuples are important, because they will be those that will carry the aggregate results, as they will be able to “see” all previous values of their groupgroup before being extracted. Thus, these bits will also be used as valid-bits for the reverse butterfly network, to denote a valid tuple, also containing the summarised information after the end of the pipeline.
Step (c) Prefix-scans
@@
the entities named $n$ are distributed inside the adapted prefix scan to calculate the permutation index of the tuples to achieve the round-robin effect by the reverse butterfly network, as with the PRRA [4]. At the same time they also operate on the keykey fields, which later becomes the resultresult field containing the aggregation summary per group. For example, if the query function is to find the sum of all the keys for each of the group IDs, the entities $n$ not only perform a cumulative sum for the reverse butterfly indexindex, but also for the key sum per group, until the last entry of the same group is processed to carry the full sum of the group. An equality check on the group field with the previous entry ($group_{i}\stackrel{{\scriptstyle?}}{{=}}group_{i-1}$) is enough for knowing where not to propagate the previous result. Using sum as an example, the start of a new groupgroup would leave the current keykey intact instead of adding the result of the last tuple.
@@
In order to also support the distinct count operator (numerical example illustrated in figure 16, right), the entities $n$ also need to follow a simple distributed algorithm: if the propagated minimum from the input node of the larger key differs from the smaller key, then the distinct counts are coming from disjoint sets, hence $n$ adopts their sum. If the minimum is equal to the smaller key of the two, then $n$ adopts this sum of distinct counts, but subtracts 1 for taking care of the double counting of the common key.
Step (d) Rollover and summarisation
@@
the entities $n^{\prime}$ are special versions of the entities $n$ and their main purpose is to output the originally-marked tuples as valid to the permutation network, while replacing their key field with the result. The main difference with $n$ is that sometimes the partial result information cannot be contained in the keykey field, and they behave as entrypoints for taking additional signals into account for these final results.
@@
For instance, if function_selectfunction\_select indicates that the operation is to find the average values per group, the entities $n$ operate the same as in the sum operation. However, it is the entity $n^{\prime}$ that will divide the result by the corresponding group tuple count (cardinality). In such a case, the entities $n$ also include internal signals for counting the number of tuples in the group in a similar fashion to the sum operation, but adding 1 instead of the key. A related difference is with the count signals of $n^{\prime}$, since $n^{\prime}$ are responsible for the rolling effect, the propagated count of the bottom $n^{\prime}$ is able to count beyond $P$ elements, e.g. by having a 32-bit width. All $n^{\prime}$ elements use this value plus the $(log_{2}P+1)$-bit wide count from the prefix sum to provide the final 32-bit count in the result, when it is requested such as from a count query. According to the complexity of the supported operators, additional signals may be carried in the network, such as the minimum, that becomes of use when calculating the distinct count per group ID.
Step (e) Round-robin compaction
@@
finally, the reverse butterfly network permutes the batch of results to achieve the round-robin effect in parallel (see figure 16, left), and is inline with [4]. The reverse butterfly network is used as an optimal-complexity switch, though other permutation networks can also be used. The lastlast flag is now renamed to a validvalid bit, as it now indicates that a port has a valid result.
II-C Numerical example
Figure 2 introduces a colour-coded numerical example on how a single batch of $P$ tuples is processed through the pipeline for the sum operator, and $P=4$. As input, the 4 tuples have group IDs ranging from {a,b,c}, and they are sorted. Whether last tuples-per-group exist in the current batch is known for the first P1=3P-1=3 tuples.
@@ II-C Numerical example
Figure 2: Colour-coded numerical example: finding the average per group. The entries in the colour legend appear in the order of usage within each tuple.
See https://philippos.info/groupby_visual for an animated version online.
Figure 3: High-throughput adaptable sliding window aggregation using a six-step pipeline.
@@
In this instance, (a,2) and (b,4) are marked with T for true as the last elements with that group ID (the example assumes that the next batch also starts with c, hence (c,9) becomes (c,9,F) on the next pipeline stage. Then, the marked tuples pass through the prefix scan, which appends to them (incrementally): a cumulative sum on their values (to facilitate sum-related operations), a cumulative sum on their valid bits (to calculate the number of tuples per group), and a cumulative sum on their recently-marked last bits (that will be used as an index to the reverse butterfly network for the round-robin effect). The first two sums are in the scope of their corresponding group IDs, and it is achieved by simply comparing their group IDs before interacting with each other according to the prefix scan topology. The next step is the additional stage in the prefix scan, where the $n^{\prime}$ entities finalise the results and also update the counts according to the previous batch. Here, (a,2) and (b,4), being selected as the tuples to carry the aggregated information for their groups a and b respectively, they replace their keys with the average values per group (e.g. by dividing 12 by 2 for b). The permutation indices are not altered in this case, remaining at 0 and 1 respectively, assuming this was the first batch, or that the previous batch has outputted its last aggregate result on index P1P-1.
III Sliding window aggregation
The design from section II-B is adapted to perform sliding-window aggregation (SWAG) queries. In terms of functionality, the main difference from the base pipeline is that the input is processed in a windowed context.
Figure 4: SWAG illustration for the sum operator.
SWAG queries are usually of the form “calculate the median of the last WS elements every WA tuples” [7], where median can be replaced by any supported aggregate operator, WS is the window size, and WA is the distance between the sampled windows. Figure 4 shows an instance of the average operator, where on every 2 tuples (WA) we find the average key within the last 4 tuples (WS). In the literature, WS is also referred to as the window range, and WA is also referred to as the bulk size, slide or stride. In some contexts, there is a group ID associated with each key, thus, the engine can also provide per group information, as with the base pipeline used for group-by aggregation.
III-A Supported functionality
The proposed SWAG engine supports the operators of the base pipeline (see section II-A), plus selection-based operators including the median. Based on the query classification from [1], the supported types are Sum-like, Max-like and Sketch-like.
In contrast to software algorithms that calculate SWAG incrementally [1], which implement the window eviction post calculation [13], here, there is no requirement for an inverse function. This is because each window is processed in its entirety, which contributes to the adaptability of the SWAG engine for operators that cannot be calculated incrementally. Section VI elaborates on the applicability of the SWAG engine to other prominent aggregation workloads.
III-B Design
The input data are seamlessly sorted on-the-fly using a small-scale sorter (see section III-D). As opposed to the base pipeline, the input is not expected to be sorted, since the latency of doing so is shown to be negligible for the window sizes of interest. Figure 3 introduces the pipeline of the SWAG engine in steps, for data arriving in batches of $P$ tuples.
Step (a) Window generation
starting from the leftmost component of figure 3, a simple buffering arrangement passes the entire window to the second component of the pipeline, which is the sorter. The buffering is used to produce WS-sized chunks (i.e. windows) while advancing by WA without the need to “replay” the incoming stream from an external module. This component uses a block ram-based circular buffer of size 2WS. Having the ability to store two separate windows is crucial for being able to seek back up to a window size (when window advance (WA) is at its highest possible value (WS)), as well as to efficiently support long bursts and backpressure in real systems. Due to the nature of the SWAG queries, the output rate has the ability to surpass the input rate, such as when WA<<WS and every element needs to produce a summary (e.g. group IDs are always unique). Hence, a “ready” signal is used to enforce backpressure to the input source.
Step (b) Sorting and cardinality count